# Interplay between local and global order controls bifurcations in nonreciprocal systems without a conservation law

### v2

@@ Abstract
Similarly to the BG approximation, the resulting evolution equations in the monomer approximation are governed by Eqs. (5), but the time-dependent probabilities are different and read (the derivation is given in [55])
$\mathcal{P}^{a}_{l,n}(t)=\frac{2\mathcal{C}^{z}_{l}(1+m^{a}(t))^{l}(1+m^{b}(t))^{n}}{(1-m^{a}(t))^{l-z}(1-m^{b}(t))^{n-1}},$ (B2)
@@
and $\mathcal{P}^{b}_{l,n}(t)$ is obtained by replacing $m^{a}(t)$ with $m^{b}(t)$ in Eq. (B2). Since $\mathcal{P}^{a}_{l,n}(t)$ is independent of the local order $q^{\mu\nu}(t)$, this implies that $q^{\mu\nu}(t)$ is slaved by $m^{\mu}(t)$, however, $q^{\mu\nu}(t)\neq m^{\mu}(t)m^{\nu}(t)$. Performing a linear stability analysis around the trivial steady-state $m^{\mu}_{s}=0$ for $J_{a}=J_{b}=J$ and $K_{a}=-K_{b}=K$, we obtain the linear stability equation $\tau d\delta\mathbf{m}(t)/dt=\mathbf{\hat{M}}(J,K)\delta\mathbf{m}(t)$ where the elements of $\mathbf{\hat{M}}(J,K)$ are given by
$\hat{M}_{aa}$ $=\hat{M}_{bb}=2\sum_{l=0}^{z}\sum_{n=0}^{1}(2l-z)\mathcal{C}^{z}_{l}\tanh{(U^{a}_{l,n})}-1,$ (B3)
$\hat{M}_{ab}$ $=-\hat{M}_{ba}=2\sum_{l=0}^{z}\sum_{n=0}^{1}(2n-1)\mathcal{C}^{z}_{l}\tanh{(U^{a}_{l,n})}.$ (B4)
@@
Interestingly, Eq. (B3) can be expressed in terms of the steady-state value of the local order in the monomer approximation, i.e., $\hat{M}_{aa}(J,K)=zq_{s}(J,K)-1$. Therefore, the eigenvalues of $\mathbf{\hat{M}}(J,K)$ are neatly given by
$\hat{\lambda}_{\pm}(J,K)=(zq_{s}(J,K)-1)\pm i\hat{M}_{ab}(J,K).$ (B5)
@@
Herefrom it follows that the Hopf bifurcation occurs at $q_{s}=1/z$ and $K\neq 0$, such that ${\rm Re}(\hat{\lambda}_{\pm})=0$ and ${\rm Im}(\hat{\lambda}_{\pm})\neq 0$. Hence, as within the BG approximation, we also find the existence of a critical local order in the monomer approximation. Contrary to the MF approximation, the Hopf line is not a straight vertical line in the $(J,K)$-plane, as shown in Fig. A1b.
@@
Appendix C: Local and global order in various approximations.—To provide a concise overview of the various approximation techniques, we summarize in Table 1 the respective evolution and conditions for the Hopf bifurcation.
Supplementary Material for: Interplay between local and global order controls bifurcations in nonreciprocal systems without a conservation law
@@
In this Supplemental Material, we provide derivations and further details about the results shown in the Letter. Furthermore, in Sec. S6 we prove that static global order in the square-lattice Ising model persists at least under small perturbations in the nonreciprocal coupling $0<|K|\leq\mathcal{O}(1/N)$ for any finite but arbitrarily large system size $N\in[m,\infty)$ for some sufficiently large $m$.
Appendix S1 Setup: Dual lattice topology and interactions
Consider a pair of lattices, denoted by $\mu=a,b$, each having a coordination number $z$ and periodic boundary conditions. On each lattice there are $N$ spins that can be in two states $\sigma^{\mu}_{i}=\pm 1$, with $i\in\{1,...,N\}$ enumerating the spin’s location. Each spin interacts with its $z$ nearest-neighbors on the same lattice and one spin at the equivalent position on the opposing lattice. The local interaction energy of spin $i$ on lattice $\mu$ can be written as (see also [1])
$E^{\mu}_{i}=-J_{\mu}\sigma^{\mu}_{i}\sum\nolimits_{\langle i|j\rangle}\sigma^{\mu}_{j}-K_{\mu}\sigma^{a}_{i}\sigma^{b}_{i},$ (S1)
where $\langle i|j\rangle$ denotes a sum over nearest-neighbors $j$ of spin $i$. Throughout, we will express energies in units of $k_{\rm B}T$. The parameters $J_{\mu}$ and $K_{\mu}$ denote the coupling within one and between the two lattices, respectively. When $K_{a}\neq K_{b}$, equivalent spins in the opposing lattices interact nonreciprocally.
Appendix S2 Glauber dynamics
We now specify the dynamics. We consider single spin-flip dynamics according to Glauber [2]. Let $P(\boldsymbol{\sigma};t)$ be the probability at time $t$ to be in state $\boldsymbol{\sigma}=\{\sigma_{1}^{a},\sigma_{1}^{b},...,\sigma_{N}^{a},\sigma_{N}^{b}\}$. This state can be reached by a single spin flip from the $2N$ distinct states $\boldsymbol{\sigma}^{\prime}_{\mu,i}=\{\sigma_{1}^{a},\sigma_{1}^{b},...,-\sigma^{\mu}_{i},...,\sigma_{N}^{a},\sigma_{N}^{b}\}$, which enters the master equation as follows
$\frac{{\rm d}P(\boldsymbol{\sigma};t)}{{\rm d}t}=\sum_{\mu=a,b}\sum_{i=1}^{N}[w^{\mu}_{i}(-\sigma^{\mu}_{i})P(\boldsymbol{\sigma}^{\prime}_{\mu,i};t)-w^{\mu}_{i}(\sigma^{\mu}_{i})P(\boldsymbol{\sigma};t)].$ (S2)
The respective transition rates $w^{\mu}_{i}(\sigma^{\mu}_{i})$ to flip a single spin $\sigma^{\mu}_{i}\rightarrow-\sigma^{\mu}_{i}$ can be specified uniquely by limiting the interactions to nearest neighbors, imposing isotropy in position space, and requiring that for $K_{a}=K_{b}$ the transition rates obey detailed-balance. These physical restrictions then lead to the general result
$w^{\mu}_{i}(\sigma^{\mu}_{i})=(1/2\tau)\left[1-\tanh{\left(\Delta E^{\mu}_{i}/2\right)}\right],$ (S3)
where $\Delta E^{\mu}_{i}$ is the change in energy on the $\mu=a\lor b$ lattice after flipping spin $\sigma^{\mu}_{i}\rightarrow-\sigma^{\mu}_{i}$, and $\tau$ is an intrinsic timescale to attempt a single spin-flip. In a discrete-time setting the intrinsic flip rate $1/\tau$ is replaced by the probability to flip a single spin, which is $1/2N$ (see also Fig. S2). According to Eq. (S1), the change in energy after such a spin-flip is given by
$\Delta E^{\mu}_{i}=-2E^{\mu}_{i}.$ (S4)
From the master equation (S2), we can directly derive equations for the first two moments of single-spin values (see also Eqs. (28) and (29) in [2])
$\tau\frac{{\rm d}\langle\sigma^{\mu}_{i}(t)\rangle}{{\rm d}t}+\langle\sigma^{\mu}_{i}(t)\rangle$ $=\langle\sigma^{\mu}_{i}(t)\tanh{(\Delta E^{\mu}_{i}/2)}\rangle,$ (S5)
$\tau\frac{{\rm d}\langle\sigma^{\mu}_{i}(t)\sigma^{\nu}_{j}(t)\rangle}{{\rm d}t}+2\langle\sigma^{\mu}_{i}(t)\sigma^{\nu}_{j}(t)\rangle$ $=\langle\sigma^{\mu}_{i}(t)\sigma^{\nu}_{j}(t)[\tanh{(\Delta E^{\mu}_{i}/2)}+\tanh{(\Delta E^{\nu}_{j}/2)}]\rangle,$ (S6)
where $\langle f(t)\rangle\equiv\sum_{\boldsymbol{\sigma}}P(\boldsymbol{\sigma};t)f(\boldsymbol{\sigma})$. Eqs. (S5)-(S6) are exact (but not yet closed) equations for the first two moments evolving under Glauber dynamics, and will serve as our starting point to derive equations for the global and local order parameters.
Appendix S3 Differential equations for the global and local order
Recall from the Letter that the global and local order are defined as
$m^{\mu}(t)$ $\equiv\frac{1}{N}\sum_{i=1}^{N}\langle\sigma^{\mu}_{i}(t)\rangle,$ (S7)
$q^{\mu\mu}(t)$ $\equiv\frac{2}{zN}\sum_{i=1}^{N}\sum_{\langle i|j\rangle}\langle\sigma_{i}^{\mu}(t)\sigma_{j}^{\mu}(t)\rangle,$ (S8)
$q^{ab}(t)$ $\equiv\frac{1}{N}\sum_{i=1}^{N}\langle\sigma_{i}^{a}(t)\sigma_{i}^{b}(t)\rangle.$ (S9)
Here, we rewrite Eqs. (S5)-(S6) in terms of differential equations for (S7)-(S9), which we do in two steps:
  1. 1.
    First, we sum Eqs. (S5)-(S6) over all spins and spin pairs. Upon doing this, the l.h.s. of Eqs. (S5)-(S6) directly transforms into
    $\frac{1}{N}\sum_{i=1}^{N}\left(\tau\frac{{\rm d}\langle\sigma^{\mu}_{i}(t)\rangle}{{\rm d}t}+\langle\sigma^{\mu}_{i}(t)\rangle\right)\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ =\tau\frac{{\rm d}m^{\mu}(t)}{{\rm d}t}+m^{\mu}(t),$ (S10)
    $\frac{2}{zN}\sum_{i=1}^{N}\sum_{\langle i|j\rangle}\left(\tau\frac{{\rm d}\langle\sigma^{\mu}_{i}(t)\sigma^{\mu}_{j}(t)\rangle}{{\rm d}t}+2\langle\sigma^{\mu}_{i}(t)\sigma^{\mu}_{j}(t)\rangle\right)=\tau\frac{{\rm d}q^{\mu\mu}(t)}{{\rm d}t}+2q^{\mu\mu}(t),$ (S11)
    $\frac{1}{N}\sum_{i=1}^{N}\left(\tau\frac{{\rm d}\langle\sigma^{a}_{i}(t)\sigma^{b}_{i}(t)\rangle}{{\rm d}t}+2\langle\sigma^{a}_{i}(t)\sigma^{b}_{i}(t)\rangle\right)\ \ \ \ \ \ \ \ =\tau\frac{{\rm d}q^{ab}(t)}{{\rm d}t}+2q^{ab}(t).$ (S12)
  2. 2.
    Next, we need to evaluate the r.h.s. of Eqs. (S5)-(S6) after summation over all spins and spin pairs. To do this, we make use of the fact that $\Delta E^{\mu}_{i}$ can take on a discrete (enumerable) set of values. Let us consider a spin with $l\in\{0,1,...,z\}$ neighboring up spins on the same lattice, and $n\in\{0,1\}$ neighboring up spins on the opposing lattice. We want to compute the change in energy upon flipping this spin. Based on Eq. (S4) we can parameterize the change in energy after flipping a spin with such an environment as follows,
    $\Delta E^{\mu}_{i}=2\sigma^{\mu}_{i}U^{\mu}_{l,n},$ (S13)
    where
    $U^{\mu}_{l,n}=[2l-z]J_{\mu}+[2n-1]K_{\mu}.$ (S14)
    This gives a direct parameterization of the energy change in terms of the number of neighboring up spins on the same $(l)$ and opposing $(n)$ lattice. Using this parameterization, we can evaluate the r.h.s. of Eqs. (S5)-(S6) using the following identities
    $\!\!\!\frac{1}{N}\sum_{i=1}^{N}\langle\sigma^{\mu}_{i}\tanh{(\Delta E^{\mu}_{i}/2)}\rangle$ $=\frac{1}{N}\sum_{i=1}^{N}\langle\tanh{(U^{\mu}_{l,n})}\rangle=\sum_{l=0}^{z}\sum_{n=0}^{1}\mathcal{P}^{\mu}_{l,n}(t)\tanh{(U^{\mu}_{l,n})},$ (S15)
    $\!\!\!\frac{2}{zN}\sum_{i=1}^{N}\sum_{\langle i|j\rangle}\langle\sigma^{\mu}_{i}\sigma^{\mu}_{j}\tanh{(\Delta E^{\mu}_{i}/2)}\rangle$ $=\frac{2}{zN}\sum_{i=1}^{N}\sum_{\langle i|j\rangle}\langle\sigma^{\mu}_{j}\tanh{(U^{\mu}_{l,n})}\rangle=\frac{2}{z}\sum_{l=0}^{z}\sum_{n=0}^{1}(2l-z)\mathcal{P}^{\mu}_{l,n}(t)\tanh{(U^{\mu}_{l,n})},$ (S16)
    $\!\!\!\frac{1}{N}\sum_{i=1}^{N}\langle\sigma^{a}_{i}\sigma^{b}_{i}\tanh{(\Delta E^{a}_{i}/2)}\rangle$ $=\frac{1}{N}\sum_{i=1}^{N}\langle\sigma^{b}_{i}\tanh{(U^{a}_{l,n})}\rangle=\sum_{l=0}^{z}\sum_{n=0}^{1}(2n-1)\mathcal{P}^{\mu}_{l,n}(t)\tanh{(U^{a}_{l,n})}.$ (S17)
For the first equality in Eqs. (S15)-(S17) we used that $\tanh{(\Delta E^{\mu}_{i}/2)}=\tanh{(\sigma^{\mu}_{i}U^{\mu}_{l,n})}=\sigma^{\mu}_{i}\tanh{(U^{\mu}_{l,n})}$, together with $(\sigma^{\mu}_{i})^{2}=1$. For the second equality, we used that terms like $\langle\tanh{(U^{\mu}_{l,n})}\rangle$ represent a weighted sum over all possible combinations of the possible values that $\tanh{(U^{\mu}_{l,n})}$ can attain. The weights are given by the time-dependent probability $\mathcal{P}^{\mu}_{l,n}(t)$ to find an up or down spin with a specific local environment. By definition, this probability is normalized as,
$\sum_{l=0}^{z}\sum_{n=0}^{1}\mathcal{P}^{\mu}_{l,n}(t)=1.$ (S18)
Combining Eqs. (S10)-(S12) and (S15)-(S17) we obtain
$\tau\frac{dm^{\mu}(t)}{dt}+m^{\mu}(t)$ $=\sum_{l=0}^{z}\sum_{n=0}^{1}\mathcal{P}^{\mu}_{l,n}(t)\tanh(U^{\mu}_{l,n}),$ (S19)
$\tau\frac{dq^{\mu\mu}(t)}{dt}+2q^{\mu\mu}(t)$ $=\frac{2}{z}\sum_{l=0}^{z}\sum_{n=0}^{1}(2l-z)\mathcal{P}^{\mu}_{l,n}(t)\tanh(U^{\mu}_{l,n}),$ (S20)
$\tau\frac{dq^{ab}(t)}{dt}+2q^{ab}(t)$ $=\sum_{\mu=a,b}\sum_{l=0}^{z}\sum_{n=0}^{1}(2n{-}1)\mathcal{P}^{\mu}_{l,n}(t)\tanh(U^{\mu}_{l,n}),$ (S21)
corresponding to Eq. (5) in the Letter. Our next aim is to obtain closed-form expressions for $\mathcal{P}^{\mu}_{l,n}(t)$, for which we consider two approximation techniques: the “monomer” approximation, and the Bethe-Guggenheim (BG) approximation.
S3.1 The “monomer” approximation
@@
We start with the lowest-order approximation, where we neglect nearest-neighbor correlations in the probability $\mathcal{P}^{\mu}_{l,n}$. Note, however, that already this level of approximation goes beyond the mean-field approach in [1]. Suppose we want to know the probability of finding an up or down spin on the $a$ lattice which has $l$ nearest up neighbors on the $a$ lattice and $n$ nearest up neighbors on the $b$ lattice. On the monomer level we assume perfect mixing between the up and down spins, resulting in
@@
where $N^{\mu}_{+}$ is the total number of up spins on the $\mu$ lattice, and similarly $N^{\mu}_{-}$ the total number of down spins. The same reasoning applies to the probability $\mathcal{P}^{b}_{l,n}$. Note that $N^{\mu}_{\pm}$ is generally time-dependent, but for simplicity we omit the subscript $t$. To relate $N^{\mu}_{\pm}$ to the global order we use the following relations (omitting the time dependence subscript $t$)
@@
Hence, in the monomer approximation, we find that the probability $\mathcal{P}^{\mu}_{l,n}(t)$ only explicitly depends on the global order $\mathbf{m}(t)$, and not on the local order $\mathbf{q}(t)\equiv(q^{aa}(t),q^{bb}(t),q^{ab}(t))$. Upon inserting Eq. (S27) into Eqs. (S19)-(S21) we obtain a closed set of five non-linear differential equations, where, notably, the dynamical equations for $\mathbf{q}(t)$ are slaved by the solutions for $\mathbf{m}(t)$, as the r.h.s. of Eqs. (S19)-(S21) do not depend on $\mathbf{q}(t)$.
S3.1.1 Difference with respect to mean-field approximation
Finally, let us point out the crucial difference between the mean-field on the one hand and the monomer and BG approximation on the other hand, which lies in the treatment of the averaging of the term $\langle\tanh{\left(\Delta E^{\mu}_{i}/2\right)}\rangle$. Whereas the mean-field approximation (uncontrollably) moves the average to the argument, i.e., $\langle\tanh{\left(\Delta E^{\mu}_{i}/2\right)}\rangle\approx\tanh{\left(\langle\Delta E^{\mu}_{i}/2\rangle\right)}$, the monomer and BG approximations use the fact that the value of $\Delta E^{\mu}_{i}/2$ lies in an enumerable set given by $U^{\mu}_{l,n}\equiv[2l-z]J_{\mu}+[2n-1]K_{\mu}$ with $l\in\{0,..,z\}$ and $n\in\{0,1\}$. This allows for an explicit summation
@@
S3.2 The Bethe-Guggenheim “pair” approximation
In the previous method, we did not take into account nearest-neighbor correlations in the evaluation of $\mathcal{P}^{\mu}_{l,n}(t)$, which we aim to include in this Section. Let us focus on the probability of picking a spin on the $a$ lattice with a given specific local environment. The same reasoning will also apply for picking a spin on the $b$ lattice. First, we split the probability in a contribution of picking an up or down spin
$\mathcal{P}^{a}_{l,n}(t)=\mathcal{P}^{a+}_{l,n}(t)+\mathcal{P}^{a-}_{l,n}(t),$ (S31)
where, e.g., $\mathcal{P}^{a+}_{l,n}(t)$ is the probability of picking an up spin with $l$ neighboring up spins on the $a$ lattice and $n$ neighboring up spins on the $b$ lattice at time $t$. On the BG level we assume ideal mixing of nearest-neighbor spin pairs, resulting in the following expression
$\mathcal{P}^{a+}_{l,n}$ $=\underbrace{[N^{a}_{+}/N]}_{\begin{subarray}{c}\rm probability\\ \begin{subarray}{c}\rm for\ up\ spin\\ {\rm on\ the}\ a\ {\rm lattice}\end{subarray}\end{subarray}}\times\underbrace{\left[\binom{N^{aa}_{++}}{l}\binom{N^{aa}_{+-}/2}{z-l}/\binom{N^{aa}_{++}+N^{aa}_{+-}/2}{z}\right]}_{\begin{subarray}{c}\rm probability\ for\\ \begin{subarray}{c}l\ {\rm\ neighboring\ up}\\ {\rm spins\ on\ the}\ a\ {\rm lattice}\end{subarray}\end{subarray}}\times\underbrace{\left[\binom{N^{ab}_{++}}{n}\binom{N^{ab}_{+-}}{1-n}/\binom{N^{ab}_{++}+N^{ab}_{+-}}{1}\right]}_{\begin{subarray}{c}\rm probability\ for\\ \begin{subarray}{c}n\ {\rm\ neighboring\ up}\\ {\rm spins\ on\ the}\ b\ {\rm lattice}\end{subarray}\end{subarray}},$ (S32)
$\mathcal{P}^{a-}_{l,n}$ $=\underbrace{[N^{a}_{-}/N]}_{\begin{subarray}{c}\rm probability\\ \begin{subarray}{c}\rm for\ down\ spin\\ {\rm on\ the}\ a\ {\rm lattice}\end{subarray}\end{subarray}}\times\underbrace{\left[\binom{N^{aa}_{-+}/2}{l}\binom{N^{aa}_{--}}{z-l}/\binom{N^{aa}_{-+}/2+N^{aa}_{--}}{z}\right]}_{\begin{subarray}{c}\rm probability\ for\\ \begin{subarray}{c}l\ {\rm\ neighboring\ up}\\ {\rm spins\ on\ the}\ a\ {\rm lattice}\end{subarray}\end{subarray}}\times\underbrace{\left[\binom{N^{ab}_{-+}}{n}\binom{N^{ab}_{--}}{1-n}/\binom{N^{ab}_{-+}+N^{ab}_{--}}{1}\right]}_{\begin{subarray}{c}\rm probability\ for\\ \begin{subarray}{c}n\ {\rm\ neighboring\ up}\\ {\rm spins\ on\ the}\ b\ {\rm lattice}\end{subarray}\end{subarray}},$ (S33)
$2N^{\mu\mu}_{\pm\pm}+N^{\mu\mu}_{+-}$ $=zN^{\mu}_{\pm},$ (S34)
$N^{ab}_{\pm\pm}+N^{ab}_{\pm\mp}$ $=N^{a}_{\pm},$ (S35)
in combination with Eq. (S25). Furthermore, we use the definition of local order (see Eqs. (S8)-(S9)) to write
$q^{\mu\mu}$ $=2(N^{\mu\mu}_{++}+N^{\mu\mu}_{--}-N^{\mu\mu}_{+-})/zN=1-4N^{\mu\mu}_{+-}/zN,$ (S36)
$q^{ab}$ $=(N^{ab}_{++}+N^{ab}_{--}-N^{ab}_{+-}-N^{ab}_{-+})/N$
$=1-2N^{ab}_{+-}/N-2N^{ab}_{-+}/N$
$=1+m^{a}-m^{b}-4N^{ab}_{+-}/N,$ (S37)
where in the last line we used the exact relation
$m^{b}-m^{a}=2(N^{ab}_{-+}-N^{ab}_{+-})/N.$ (S38)
Using the relations (S34)-(S37) we obtain the following result for the spin pairs within the same lattice
$N^{\mu\mu}_{\pm\pm}$ $=(z/8)N(1\pm 2m^{\mu}+q^{\mu\mu}),$ (S39)
$N^{\mu\mu}_{\pm\mp}$ $=(z/4)N(1-q^{\mu\mu}),$ (S40)
and for the spin pairs between the two opposing lattices
$N^{ab}_{\pm\pm}$ $=(1/4)N(1\pm m^{a}\pm m^{b}+q^{ab}),$ (S41)
$N^{ab}_{\pm\mp}$ $=(1/4)N(1\pm m^{a}\mp m^{b}-q^{ab}).$ (S42)
Inserting these relations back into Eqs. (S32)-(S33) and taking the thermodynamic limit $N\rightarrow\infty$ while keeping $\mathbf{m}(t)$ and $\mathbf{q}(t)$ fixed, we obtain (omitting, for convenience, arguments on the l.h.s.)
$\mathcal{P}^{a+}_{l,n}(t)$ $\simeq\frac{\mathcal{C}^{z}_{l}(1+2m^{a}(t)+q^{aa}(t))^{l}(1+m^{a}(t)-m^{b}(t)-q^{ab}(t))^{1-n}}{(1+m^{a}(t))^{z}(1-q^{aa}(t))^{l-z}(1+m^{a}(t)+m^{b}(t)+q^{ab}(t))^{-n}},$ (S43)
$\mathcal{P}^{a-}_{l,n}(t)$ $\simeq\frac{\mathcal{C}^{z}_{l}(1-2m^{a}(t)+q^{aa}(t))^{z-l}(1-m^{a}(t)-m^{b}(t)+q^{ab}(t))^{1-n}}{(1-m^{a}(t))^{z}(1-q^{aa}(t))^{-l}(1-m^{a}(t)+m^{b}(t)-q^{ab}(t))^{-n}},$ (S44)
where $\mathcal{C}^{z}_{l}$ is given in Eq. (S29). Expressions (S43)-(S44) are given in the main Letter in Eq. (6). The expressions for $\mathcal{P}^{b\pm}_{l,n}(t)$ are similar to Eqs. (S43)-(S44) upon interchanging $m^{a}(t)$ and $m^{b}(t)$ and replacing $q^{aa}(t)$ with $q^{bb}(t)$. The difference with the “monomer” approximation (S27) and (S28) is that in the BG approximation the probability also depends explicitly on the local order $\mathbf{q}(t)$. Inserting Eqs. (S43)-(S44) into Eq. (S31), and finally into Eqs. (S19)-(S21), leads to a closure of the five coupled non-linear differential equations.

### v3

@@ Appendix F Monomer approximation
Figure G1: Eigenvalues $l_{k}$ of the Laplacian matrix $\mathbf{L}$ for the mean-field lattice (a), Bethe lattice (b), and square lattice (c). For the mean-field lattice, all eigenvalues except for the first are degenerate with value $l_{2}=...=l_{N}=N$, where $N$ is the system size given in table 2. The Bethe lattice has a spectral gap between the lowest eigenvalue $l_{1}=0$ and $l_{2}$.
@@
Similar to the BG approximation, the resulting evolution equations in the monomer approximation are governed by Eqs. (8)-(10), but the time-dependent probabilities are different and read (the derivation is given in Note (2))
$\mathcal{P}^{a}_{l,n}(t)=\frac{2\mathcal{C}^{z}_{l}(1+m^{a}(t))^{l}(1+m^{b}(t))^{n}}{(1-m^{a}(t))^{l-z}(1-m^{b}(t))^{n-1}},$ (F2)
@@
and $\mathcal{P}^{b}_{l,n}(t)$ is obtained by replacing $m^{a}(t)$ with $m^{b}(t)$ in Eq. (F2). Since $\mathcal{P}^{a}_{l,n}(t)$ is independent of the local order $q^{\mu\nu}(t)$, this implies that $q^{\mu\nu}(t)$ is slaved by $m^{\mu}(t)$, however, $q^{\mu\nu}(t)\neq m^{\mu}(t)m^{\nu}(t)$. Performing a linear stability analysis around the trivial steady state $m^{\mu}_{s}=0$ for $J_{a}=J_{b}=J$ and $K_{a}=-K_{b}=K$, we obtain a linear stability equation where the eigenvalues of the linear stability matrix can neatly be written as
$\hat{\lambda}_{\pm}(q_{s};J,K)=(zq_{s}(J,K)-1)\pm i\hat{M}_{2}(J,K),$ (F3)
@@
where $q_{s}(J,K)$ is the steady-state value of the local order in the monomer approximation, and
$\hat{M}_{2}(J,K)=2\sum_{l=0}^{z}\sum_{n=0}^{1}(2n-1)\mathcal{C}^{z}_{l}\tanh{(U^{a}_{l,n})}.$ (F4)
@@
From this follows that the Hopf bifurcation occurs at $q_{s}=1/z$ and $K\neq 0$, such that ${\rm Re}(\hat{\lambda}_{\pm})=0$ and ${\rm Im}(\hat{\lambda}_{\pm})\neq 0$. Hence, as within the BG approximation, we also find the existence of a critical local order in the monomer approximation. Contrary to the MF approximation, the Hopf line is not a straight vertical line in the $(J,K)$-plane, as shown in Fig. E1b.
@@
To provide a concise overview of the various approximation techniques, we summarize in Table 1 the respective evolution and conditions for the Hopf bifurcation.
Appendix G Kinetic Monte-Carlo simulations
For the results shown in Fig. 1c,d (black dashed lines) and Fig. 4 we performed kinetic MC simulations on three different types of lattices: (i) the fully connected mean-field (MF) lattice, (ii) the Bethe lattice, and (iii) the square lattice with periodic boundary conditions. Simulations on the Bethe lattice were performed using the random graph algorithm Johnston and Plechác (1998); Dhar et al. (1997), which works as follows: Consider a Bethe lattice with coordination number $z$. First, we create a Cayley tree of $i=\{1,...,N\}$ spins with coordination number $z$. The spins on the outer layer are connected to one spin on the inner layer. To create the remaining $z-1$ connections, we randomly pair spins on the outer layer to other spins on the outer layer. The final result is a Cayley tree with random connections on the outer layer. Note that for both lattices $a$ and $b$ we create new random connections. For large $N$, it has been shown that the Ising model on an ensemble of such random graphs is equivalent to the Ising model on a Bethe lattice Johnston and Plechác (1998). Indeed, for large $N$ we find perfect agreement between the simulations and our theory, as shown in Fig. 1c,d. For the MF lattice, we connect all spins with each other, resulting in a fully connected graph.
G1 Simulation setup
In Table 2 we summarize the size of the system, the number of trajectories, and the parameter settings that were used to obtain the spectral density shown in Fig. 4. As initial conditions, we selected a randomly mixed configuration of up- and down-spins for fixed magnetization.
Table 2: Simulation parameters for results shown in Fig. 4.
lattice size ($N$) MC steps # traj. $J$ $[k_{\rm B}T]$ $K$ $[k_{\rm B}T]$
mean-field $1500$ $N\times 10^{3}$ $500$ $1.5/N$ $0.3$
Bethe $1457$ $N\times 10^{3}$ $500$ $0.5$ $0.3$
square $40\times 40$ $N\times 10^{3}$ $500$ $0.6$ $0.3$
For the results shown in Fig. 1c,d we used a Bethe lattice with system size $N=118097$. Such a large system size was not feasible for the setup of Fig. 4 since the spectral density $\langle|\hat{\Psi}^{\mu}_{k}(\omega)|\rangle$ must be averaged over many independent trajectories, resulting in memory issues for too large $N$.
G2 Eigenvalues of Laplacian matrix
In Fig. 4 we plot the spectral density $\langle|\hat{\Psi}^{\mu}_{k}(\omega)|\rangle$ as a function of the eigenvalues $l_{k}$ of the Laplacian matrix $\mathbf{L}$. To obtain the eigenvalues, we numerically diagonalized the Laplacian $\mathbf{L}$ in Python, and the resulting eigenvalues are shown in Fig. G1. Note that for the mean-field lattices, all eigenvalues except the first are degenerate and equal to $l_{2}=...=l_{N}=N$.
G3 Animations
To visualize the dynamics of the nonreciprocal Ising model on the mean-field lattice, Bethe lattice, and square lattice, we have provided animations in Note (2). In each animation, we show three independent simulations on the aforementioned lattices, with the coupling strengths $(J,K)$ reported in Table 2. More information about the animations is given below:
  • “MF_lattice.gif” shows simulations on the mean-field lattice for $N=2000$ spins, where each spin is connected to every other spin. For illustrative purposes, the edges between spins are not shown.
  • “Bethe_lattice.gif” shows simulations on the Bethe lattice for $N=13121$ spins, corresponding to a Bethe lattice with $8$ layers and a coordination number of $z=4$. For illustrative purposes, only the first $6$ layers of the Bethe lattice are shown.
  • “Square_lattice.gif” shows simulations on the square lattice for $N=122\times 122$ spins.
Supplementary Material for: Local Order Controls the Onset of Oscillations in the Nonreciprocal Ising Model
@@ Appendix G Kinetic Monte-Carlo simulations
In this Supplementary Material, we provide derivations and further details of the results shown in the main paper. Furthermore, in Sec. S3 we prove that static global order in the square-lattice Ising model persists at least under small perturbations in the nonreciprocal coupling $0<|K|\leq\mathcal{O}(1/N)$ for any finite but arbitrarily large system size $N\in[m,\infty)$ for some sufficiently large $m$.
Appendix S1 The “monomer” approximation
@@
Here we introduce the monomer approximation to obtain an approximate expression for the probability $\mathcal{P}^{\mu}_{l,n}(t)$. The final results are also shown in Appendix F of the main paper. Suppose we want to know the probability of finding an up- or down-spin on the $a$ lattice that has $l$ up nearest neighbors on the $a$ lattice and $n$ nearest up neighbors on the $b$ lattice. On the monomer level we assume perfect mixing between the up and down spins, resulting in
@@
where $N^{\mu}_{+}$ is the total number of up spins on the $\mu$ lattice, and similarly $N^{\mu}_{-}$ the total number of down spins. The same reasoning applies to the probability $\mathcal{P}^{b}_{l,n}$. Note that $N^{\mu}_{\pm}$ is generally time dependent, but for simplicity we omit the explicit time dependence. To relate $N^{\mu}_{\pm}$ to the global order we use the following relations
@@
Hence, in the monomer approximation, we find that the probability $\mathcal{P}^{\mu}_{l,n}(t)$ only explicitly depends on the global order $\mathbf{m}(t)$, and not on the local order $\mathbf{q}(t)\equiv(q^{aa}(t),q^{bb}(t),q^{ab}(t))$.
S1.1 Difference with respect to mean-field approximation
Finally, let us point out the crucial difference between the mean-field approximation on the one hand and the monomer and BG approximation on the other hand, which lies in the treatment of the averaging of the term $\langle\tanh{\left(\Delta E^{\mu}_{i}/2\right)}\rangle$. Whereas the mean-field approximation (uncontrollably) moves the average to the argument, i.e. $\langle\tanh{\left(\Delta E^{\mu}_{i}/2\right)}\rangle\approx\tanh{\left(\langle\Delta E^{\mu}_{i}/2\rangle\right)}$, the monomer and BG approximations use the fact that the value of $\Delta E^{\mu}_{i}/2$ lies in an enumerable set given by $U^{\mu}_{l,n}\equiv[2l-z]J_{\mu}+[2n-1]K_{\mu}$ with $l\in\{0,..,z\}$ and $n\in\{0,1\}$. This allows for an explicit summation